The oxygen uptake for cardiac patients who regularly exercise follows a normal distribution with a mean of 24.1 ml/kg and a standard deviation of 6.30 ml/kg. What is the value of oxygen uptake of the top 10% of the population

Select one:

a. -30.90

b. 0.14

c. 30.90

d. 1.08

e. 15.45
Answer #1

Solution :

Given that ,

mean = = 24.1

standard deviation = = 6.30

The z - distribution of the 10 % is,

P( Z > z ) = 10 %

1 - P( Z < z ) = 0.10

P( Z <  ) = 1 - 0.10

P( Z < z ) = 0.90

P( Z < 1.282 ) = 0.90

z = 1.282

Using z - score formula,

X = z * +

= 1.282 * 3.60 + 24.1

= 28.7